On January 1, 2021, the Mason Manufacturing Company began construction of a building to be used as its office headquarters. TComplete this question by entering your answers in the tabs below. Reg 1 and 3 Req 2 Calculate the amount of interest that MaComplete this question by entering your answers in the tabs below. Req 1 and 3 Reg 2 What is the total cost of the building?

No explanations are needed. Just give me the correct answers! Thank you so much.
Answer #1

Working Notes

Calculation of Interest Amount
Year Particulars Loan Amount ($) Rate of Interest ($) Years Interest Amount ($)
2021 Capitalisation of Interest on Construction Loan             3,000,000 15% 1.00 450,000
2022             3,000,000 15% 0.75 337,500
Interest Expense
2021 On construction Loan 0
On Lolg term Note 1             3,000,000 9%    1.00 270,000
On Lolg term Note 1             9,000,000 13%    1.00 1,170,000
1,440,000
2022 On construction Loan             3,000,000 15% 0.25 112,500
On Lolg term Note 1             3,000,000 9% 1.00 270,000
On Lolg term Note 1             9,000,000 13% 1.00 1,170,000
1,552,500

Total Cost of Building Cost = Total expenditure on the project + Interest capitalization

Total Expenditure on the project = 1,170,000+840,000+1,040,000+840,000+306,000+639,000+900,000

= 5,735,000

Interest Capitalization = 450,000+337,500

= 787,500

Therefore Total Building Cost = 5,735,000 + 787,500

=6,522,500

Required Information are as follows

Particular 2021 2022
Interest Capitalized in ($) 450,000 337,500
Interest Expense in ($) 1,440,000 1,552,500
Total Cost of Building $ 6,522,500